# TableMaster: A Recipe to Advance Table Understanding with Language Models

### v1

@@ 1 Introduction
Tables are widely used in daily life and across various fields, such as healthcare (Ghasemi & Amyot, 2016) and finance (Li et al., 2020), due to their unique ability to efficiently represent two-dimensional relational data. It is crucial to process tabular data with both efficiency and accuracy. Recently, large language models (LLMs) (Gunasekar et al., 2023; OpenAI, 2024; Touvron et al., 2023) have achieved significant progress in the field of natural language processing. They perform well in a wide range of downstream text-based tasks, including language understanding (Minaee et al., 2024; Zhu et al., 2024) and reasoning (Plaat et al., 2024). Naturally, language models (LMs) are increasingly being used to process and understand tabular data (Fang et al., 2024; Zhang et al., 2024b), enabling reasoning for downstream tasks such as table-based question answering (Pasupat & Liang, 2015) and table-based fact verification (Chen et al., 2020).
However, the data structure of tables inherently possess a unique two-dimensional structure that contrasts with the linear text, which dominates the content in language model pretraining corpora. Most advanced LMs are not specifically optimized for processing tabular data. While techniques such as chain-of-thought prompting (Wei et al., 2023) and other reasoning-enhanced methods (Yao et al., 2023) have enabled LMs to perform satisfactorily in reasoning with linear text, significant room for improvement remains in table-based reasoning (Chen, 2023). A notable gap persists in LMs’ ability to fully understand tables and effectively reason with tabular data.
@@
Many previous studies have aimed to improve the table understanding capabilities of LMs. One efficient approach is using prompting to adapt LMs for table understanding without requiring fine-tuning, making it applicable to any advanced LM. Recent studies primarily adopt two main strategies to enhance table understanding with LMs. The first strategy involves extracting a sub-table that contains relevant content from the original table to reduce the context size, thereby making it easier for LMs to comprehend. Examples include Dater (Ye et al., 2023) and Chain-of-Table (Wang et al., 2024), among others. The second strategy leverages SQL or Python programs to augment numerical reasoning, locate target data, and enhance table understanding of numerical information, as demonstrated by Binder (Cheng et al., 2023) and LEVER (Ni et al., 2023), etc. However, these studies primarily focus on a single basic aspect to enhance the performance of LMs in table understanding or design complex methods with isolated strategies. There is currently an absence of work that provides a systematic and fundamental analysis of table understanding with language models and proposes comprehensive methods for its improvement.
@@
2 Related Work
Reasoning with Language Models. It has been observed that language models (LMs) can exhibit reasoning abilities when they are sufficiently large (Wei et al., 2022; Suzgun et al., 2022). LMs are now widely used for various reasoning tasks, such as question answering (Kamalloo et al., 2023), decision making (Yang et al., 2023), and mathematical reasoning (Ahn et al., 2024). At the inference stage, techniques such as chain-of-thought prompting (Wei et al., 2023) are used to trigger step-by-step reasoning processes and improve reasoning performance. Few-shot prompting (Brown et al., 2020), least-to-most prompting (Zhou et al., 2023), and program-of-thought prompting (Chen et al., 2023) have proven effective in specific scenarios. Methods like self-consistency (Wang et al., 2023b) and structuring the reasoning process in forms like trees (Yao et al., 2023) or graphs (Besta et al., 2024; Cao, 2024a) are also useful for more complex reasoning tasks. Recently, many works have focused on using reinforcement learning (Lightman et al., 2023; Uesato et al., 2022) to improve the reasoning abilities of LMs during training. Our work focuses on inference-time improvements and proposes a general framework applicable to all kinds of LMs for table understanding and reasoning.
Fine-Tuning LMs for Table Understanding. Several studies have focused on fine-tuning language models to enhance their understanding of tabular data. For example, based on the masked language modeling approach introduced in BERT (Devlin et al., 2019), models like TaPas (Herzig et al., 2020), Pasta (Gu et al., 2022), and TUTA (Wang et al., 2021) propose specialized pre-training methods to improve LMs’ ability to process tables. Similarly, TAPEX (Liu et al., 2022) pre-trains an encoder-decoder model to function as a SQL executor, enabling better table comprehension. Recent advancements, such as TableLlama (Zhang et al., 2024a), TableGPT (Zha et al., 2023), and StructLLM (Zhuang et al., 2024), leverage open-sourced decoder-only models like Llama (Touvron et al., 2023) to pre-train larger models optimized for various downstream table-related tasks.
Adapting LMs for Table Understanding Without Fine-Tuning. Other studies focus on adapting LMs to table-related tasks without requiring fine-tuning. For instance, Binder (Cheng et al., 2023), LEVER (Ni et al., 2023), and PoTable (Mao et al., 2024) generate SQL or Python programs, extending the capabilities of LMs to analyze tabular data. Dater (Ye et al., 2023), TabSQLify (Nahid & Rafiei, 2024a), ReAcTable (Zhang et al., 2023), TAP4LLM (Sui et al., 2024), and Tree-of-Table (Ji et al., 2024) introduce different methods to construct sub-tables, modifying the tabular context for improved understanding. Chain-of-Table (Wang et al., 2024) generalizes various table operations, dynamically generating reasoning chains to create sub-tables. MIX-SC (Liu et al., 2024b) employs table normalization and leverages self-consistency, combining results from Python agents and textual reasoning to enhance performance. SpreadsheetEncoder (Dong et al., 2024) is specifically designed to interpret tabular data within spreadsheet environments. Our work also follows this direction to focus on adapting LMs without fine-tuning. We identify key challenges in table understanding and address them through our proposed method, which can be applied to any advanced LMs.
3 Challenges in Table Understanding
As illustrated in Figure 1, we identify and analyze the challenges in table understanding with language models (LMs) through the experiments shown in Figure 2 and related discussions. Additionally, we propose targeted solutions to address these challenges. The detailed settings of the challenge analysis experiment are provided in Appendix B.
Tabular Data Characteristics. Tabular data differs from regular text, which is typically linear and sequential, due to its structured nature. Although tabular data can be represented as sequential text, it is fundamentally a two-dimensional array of cells. Each cell primarily contains text, but the cells are interconnected and share relationships with one another. Typically, cells within the same column represent the same feature or type, while cells in the same row correspond to a single data instance. Tables are highly efficient for data representation, often containing a large amount of information, making them inherently data-intensive. Moreover, the text in tables is typically concise, consisting of simple words and phrases rather than continuous sentences, leading to sparse semantic context. Lastly, tables frequently include substantial amounts of numerical data, such as dates, times, scores, and measurements, which often require specialized processing.
3.1 Difficulty in Locating Target Data
@@
When LMs encounter tabular data, they often struggle to locate the target data relevant to a given query, leading to misunderstandings. This challenge arises because tabular data is inherently data-intensive, typically containing large volumes of information. Additionally, the structured nature of tabular data makes it challenging for LMs to interpret individual cell contents within the broader context of headers and other structural information. This issue can lead to long-context hallucination (Huang et al., 2024). Moreover, LMs are prone to neglecting information in the middle of the context (Liu et al., 2024a), making it even harder to locate target data and further impairing their overall comprehension of the table. (Figure 1 - C1)
@@ 3.2 Table Semantic Deficiency
Tabular data is typically concise, with most cells containing simple words or phrases. Additionally, for each data entry in a row, some descriptive information may reside outside the row, such as in the top header or other structural elements. Understanding a cell in isolation is challenging and often requires a deeper comprehension of the structural relationships within the table. This leads to the problem of sparse semantic context, which is fundamentally different from the rich semantic context found in most data used during LMs’ pretraining (Dong et al., 2022). The semantic deficiency in tables makes it difficult for LMs to effectively understand and process tabular data. (Figure 1 - C2)
@@
To address this issue, we propose a solution where tables are first verbalized into sequential, natural text as a description and then provided to LMs alongside the original table before they directly tackle table-related tasks. It is similar to table2text (Parikh et al., 2020). This transformation enriches the semantic context, making the data more aligned with the LMs’ pretraining, thereby enhancing their ability to effectively understand and process tabular data. (Figure 1 - S2)
@@ 3.3 Numerical Inaccuracy in Textual Reasoning
Tabular data often contains numerical values, such as dates, times, scores, and other recorded numbers, and is typically intensive. However, when LMs are used to process numerical data in textual reasoning, they often face significant limitations. LMs are prone to arithmetic calculation errors, especially when dealing with large numbers. LMs are also inefficient at handling iterative processes, particularly when the number of iteration steps is large (Chen et al., 2023). (Figure 1 - C3)
@@
Symbolic methods offer a promising solution to these challenges and have been explored extensively in prior research (Cheng et al., 2023; Ni et al., 2023; Mao et al., 2024). Using symbolic tools, such as SQL or Python programs in combination with LMs, provides an effective approach to handling numerical data in tabular formats. (Figure 1 - S3)
3.4 Semantic Inflexibility in Symbolic Reasoning
Symbolic methods excel at arithmetic calculations. However, when prompting LMs to generate code for program of thought reasoning, the performance is suboptimal. Instead of truly understanding the context and generating problem-solving code, LMs often rely on memorized code from the pretraining stage (Yang et al., 2024). We refer to this limitation as semantic inflexibility. In table understanding, this challenge is exacerbated by the table’s complex structure and concise text content. In real-world scenarios, noisy tables further hinder LMs’ symbolic reasoning capabilities. Consequently, while symbolic reasoning with numerical data is highly accurate, the generated code may be incorrect due to issues in program logic or data handling, leading to errors or unintended results. (Figure 1 - C4)
To address this, we first normalize the table structure and content, ensuring that each column follows a consistent format. We then propose a solution where LMs first engage in textual reasoning before generating symbolic reasoning programs. This preliminary textual reasoning step serves as a guide for subsequent symbolic reasoning, improving alignment with the task context. Our approach can be seen as encouraging LMs to think more thoroughly before reasoning, aligning with techniques like plan-and-solve (Wang et al., 2023a). By incorporating textual reasoning as a foundation, we enhance the accuracy and contextual relevance of symbolic reasoning. As demonstrated in Figure 2(c), this method achieves a higher accuracy of 59.1% for calculation-required questions. (Figure 1 - S4)
4 TableMaster: A Recipe for Table Understanding
Based on findings in Section 3, we introduce a recipe and comprehensive framework, TableMaster, as shown in Figure 3. It integrates the propose solution proposed in Section 3 into a unified recipe for table understanding. The framework encompasses three key processes: Table Structure Understanding, Table Content Understanding, and Table Reasoning for QA.
@@ 4.3 Table Content Understanding
Studies have shown that LMs can assess whether sufficient information is available to answer a question (Cao, 2024b; Yin et al., 2023). We first prompt the LMs to estimate whether the constructed Table-of-Focus $\mathbb{T}^{F}_{a\times b}$, containing $C^{0}$, provides enough information to answer the given question $Q$. If not, additional column attributes from the candidate column set $\mathbf{C}$ are incrementally added from the ranked candidate headers until sufficient information is available or all relevant top headers have been utilized. Subsequently, a total of $a^{\prime}$ columns from $C$ are selected for further reasoning. We use re-construction to mitigate information loss during the table-of-focus construction process. The detailed re-construction algorithm can be found in Appendix H.
@@ 4.4 Table Reasoning for Question Answering
Table 1: Performance comparison between TableMaster and previous work on WikiTQ and TabFact. The values in the table represent accuracy (%). The best result is bold, the second-best result is underlined, and the improvement over the previous best result is highlighted in green. ‘-’ indicates that the result values were not reported in the related papers. Our method outperforms all other methods across both datasets and different language models.
Method WikiTQ TabFact
Text-to-SQL (Rajkumar et al., 2022) 52.90 - - 64.71 - -
End-to-End QA (Wang et al., 2024) 51.84 - - 70.45 - -
Few-Shot QA (Wang et al., 2024) 52.56 - - 71.54 - -
Chain-of-Thought (Wang et al., 2024) 53.48 - - 65.37 - -
ReAcTable (Zhang et al., 2023) 52.50 - - 74.40 - -
Binder (Cheng et al., 2023) 56.74 58.86 50.51 79.17 84.63 78.16
Dater (Ye et al., 2023) 52.81 58.33 43.53 78.01 80.98 81.57
TabSQLify (Nahid & Rafiei, 2024a) 64.70 57.02 55.78 79.50 78.75 70.70
Chain-of-Table (Wang et al., 2024) 59.94 55.60 62.22 80.20 84.24 85.62
Tree-of-Table (Ji et al., 2024) 61.11 - - 81.92 - -
PoTable (Mao et al., 2024) - 64.73 65.56 - 88.93 87.06
@@ 5.1 Settings
We conduct extensive experiments to evaluate the performance of TableMaster. Specifically, we assess its effectiveness across three different table understanding datasets: WikiTQ (Pasupat & Liang, 2015) (table-based question answering), TabFact (Chen et al., 2020) (table-based fact verification), and FetaQA (Nan et al., 2022) (table-based free-form question answering). For WikiTQ and TabFact, following previous work (Wang et al., 2024; Liu et al., 2024b), we use exact match accuracy as the evaluation metric. For FetaQA, we evaluate performance using BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004) scores. Tables are encoded in Markdown format before being input into language models, with or without addresses, depending on the specific case LABEL:ap:exampleLABEL:ap:example.
Our experiments utilize OpenAI models hosted on Microsoft Azure. Unless otherwise stated, we set the temperature to 0 to ensure stable output while keeping all other hyperparameters at their default values. The models used in our evaluation include gpt-4o (gpt-4o-0806), gpt-4o-mini (gpt-4o-mini-0718), gpt-3.5-turbo (gpt-3.5-turbo-0125), o1 (o1-preview-0912), and o1-mini (o1-mini-0912). Additionally, we evaluate our methods on Llama-3.1-70B (Llama-3.1-70B-Instruct). The exact number of parameters for several LMs (e.g., GPT, o1) has not been publicly disclosed. Most parameter counts are estimates reported to provide context for understanding model performance. For more precise information, please refer to the original or future official documentation (Abacha et al., 2025).
For comparison, we select several strong baselines, including both classic and state-of-the-art methods such as Binder (Cheng et al., 2023), Dater (Ye et al., 2023), and Chain-of-Table (Wang et al., 2024). Performance results for other methods not in this work are cited directly from their original or related papers, with sources indicated alongside the method names in the results table.
Further analysis and additional experiments on TableMaster can be found in the Appendix. The prompts used in TableMaster can be found in Appendix M, while other prompts used in this work are provided in Appendix N.

### v3

@@ 1 Introduction
Tables are widely used in daily life and across various fields, such as healthcare [1] and finance [2], due to their unique ability to efficiently represent two-dimensional relational data. It is crucial to process tabular data with both efficiency and accuracy. Recently, large language models (LLMs) [3, 4, 5] have achieved significant progress in the field of natural language processing. They perform well in a wide range of downstream text-based tasks, including language understanding [6, 7] and reasoning [8]. Naturally, language models (LMs) are increasingly being used to process and understand tabular data [9, 10], enabling reasoning for downstream tasks such as table-based question answering [11] and table-based fact verification [12].
However, the data structure of tables inherently possess a unique two-dimensional structure that contrasts with the linear text, which dominates the content in language model pretraining corpora. Most advanced LMs are not specifically optimized for processing tabular data. While techniques such as chain-of-thought prompting [13] and other reasoning-enhanced methods [14] have enabled LMs to perform satisfactorily in reasoning with linear text, significant room for improvement remains in table-based reasoning [15]. A notable gap persists in LMs’ ability to fully understand tables and effectively reason with tabular data.
@@
Many previous studies have aimed to improve the table understanding capabilities of LMs. One efficient approach is using prompting to adapt LMs for table understanding without requiring fine-tuning, making it applicable to any advanced LM. Recent studies primarily adopt two main strategies to enhance table understanding with LMs. The first strategy involves extracting a sub-table that contains relevant content from the original table to reduce the context size, thereby making it easier for LMs to comprehend. Examples include Dater [16] and Chain-of-Table [17], among others. The second strategy leverages SQL or Python programs to augment numerical reasoning, locate target data, and enhance table understanding of numerical information, as demonstrated by Binder [18] and LEVER [19], etc. However, these studies primarily focus on a single basic aspect to enhance the performance of LMs in table understanding or design complex methods with isolated strategies. There is currently an absence of work that provides a systematic and fundamental analysis of table understanding with language models and proposes comprehensive methods for its improvement.
@@
2 Related Work
Reasoning with Language Models. It has been observed that language models (LMs) can exhibit reasoning abilities when they are sufficiently large [20, 21]. LMs are now widely used for various reasoning tasks, such as question answering [22], decision making [23], and mathematical reasoning [24]. At the inference stage, techniques such as chain-of-thought prompting [13] are used to trigger step-by-step reasoning processes and improve reasoning performance. Few-shot prompting [25], least-to-most prompting [26], and program-of-thought prompting [27] have proven effective in specific scenarios. Methods like self-consistency [28] and structuring the reasoning process in forms like trees [14] or graphs [29, 30] are also useful for more complex reasoning tasks. Recently, many works have focused on using reinforcement learning [31, 32] to improve the reasoning abilities of LMs during training. Our work focuses on inference-time improvements and proposes a general framework applicable to all kinds of LMs for table understanding and reasoning.
Fine-Tuning LMs for Table Understanding. Several studies have focused on fine-tuning language models to enhance their understanding of tabular data. For example, based on the masked language modeling approach introduced in BERT [33], models like TaPas [34], Pasta [35], and TUTA [36] propose specialized pre-training methods to improve LMs’ ability to process tables. Similarly, TAPEX [37] pre-trains an encoder-decoder model to function as a SQL executor, enabling better table comprehension. Recent advancements, such as TableLlama [38], TableGPT [39], and StructLLM [40], leverage open-sourced decoder-only models like Llama [5] to pre-train larger models optimized for various downstream table-related tasks.
Adapting LMs for Table Understanding Without Fine-Tuning. Other studies focus on adapting LMs to table-related tasks without requiring fine-tuning. For instance, Binder [18], LEVER [19], and PoTable [41] generate SQL or Python programs, extending the capabilities of LMs to analyze tabular data. Dater [16], TabSQLify [42], ReAcTable [43], TAP4LLM [44], and Tree-of-Table [45] introduce different methods to construct sub-tables, modifying the tabular context for improved understanding. Chain-of-Table [17] generalizes various table operations, dynamically generating reasoning chains to create sub-tables. MIX-SC [46] employs table normalization and leverages self-consistency, combining results from Python agents and textual reasoning to enhance performance. SpreadsheetEncoder [47] is specifically designed to interpret tabular data within spreadsheet environments. Our work also follows this direction to focus on adapting LMs without fine-tuning. We identify key challenges in table understanding and address them through our proposed method, which can be applied to any advanced LMs.
3 Challenges in Table Understanding
As illustrated in Figure 1, we identify and analyze the challenges in table understanding with language models (LMs) through the experiments shown in Figure 2 and related discussions. Additionally, we propose targeted solutions to address these challenges. The detailed settings of the challenge analysis experiment are provided in Appendix D.
Tabular Characteristics. Tabular data differs from regular text, which is typically linear and sequential, due to its structured nature. Although tabular data can be represented as sequential text, it is fundamentally a two-dimensional array of cells. Each cell primarily contains text, but the cells are interconnected and share relationships with one another. Typically, cells within the same column represent the same feature or type, while cells in the same row correspond to a single data instance. Tables are highly efficient for data representation, often containing a large amount of information, making them inherently data-intensive. Moreover, the text in tables is typically concise, consisting of simple words and phrases rather than continuous sentences, leading to sparse semantic context. Lastly, tables frequently include substantial amounts of numerical data, such as dates, times, scores, and measurements, which often require specialized processing.
3.1 Difficulty in Locating Target Data
@@
When LMs encounter tabular data, they often struggle to locate the target data relevant to a given query, leading to misunderstandings. This challenge arises because tabular data is inherently data-intensive, typically containing large volumes of information. Additionally, the structured nature of tabular data makes it challenging for LMs to interpret individual cell contents within the broader context of headers and other structural information. This issue can lead to long-context hallucination [48]. Moreover, LMs are prone to neglecting information in the middle of the context [49], making it even harder to locate target data and further impairing their overall comprehension of the table. (Figure 1 - C1)
@@ 3.2 Table Semantic Deficiency
Tabular data is typically concise, with most cells containing simple words or phrases. Additionally, for each data entry in a row, some descriptive information may reside outside the row, such as in the top header or other structural elements. Understanding a cell in isolation is challenging and often requires a deeper comprehension of the structural relationships within the table. This leads to the problem of sparse semantic context, which is fundamentally different from the rich semantic context found in most data used during LMs’ pretraining [50]. The semantic deficiency in tables makes it difficult for LMs to effectively understand and process tabular data. (Figure 1 - C2)
@@
To address this issue, we propose a solution where tables are first verbalized into sequential, natural text as a description and then provided to LMs alongside the original table before they directly tackle table-related tasks. It is similar to table2text [51]. This transformation enriches the semantic context, making the data more aligned with the LMs’ pretraining, thereby enhancing their ability to effectively understand and process tabular data. (Figure 1 - S2)
@@ 3.3 Numerical Inaccuracy in Textual Reasoning
Tabular data often contains numerical values, such as dates, times, scores, and other recorded numbers, and is typically intensive. However, when LMs are used to process numerical data in textual reasoning, they often face significant limitations. LMs are prone to arithmetic calculation errors, especially when dealing with large numbers. LMs are also inefficient at handling iterative processes, particularly when the number of iteration steps is large [27]. (Figure 1 - C3)
@@
Symbolic methods offer a promising solution to these challenges and have been explored extensively in prior research [18, 19, 41]. Using symbolic tools, such as SQL or Python programs in combination with LMs, provides an effective approach to handling numerical data in tabular formats. (Figure 1 - S3)
3.4 Semantic Inflexibility in Symbolic Reasoning
Symbolic methods excel at arithmetic calculations. However, when prompting LMs to generate code for program of thought reasoning, the performance is suboptimal. Instead of truly understanding the context and generating problem-solving code, LMs often rely on memorized code from the pretraining stage [52]. We refer to this limitation as semantic inflexibility. In table understanding, this challenge is exacerbated by the table’s complex structure and concise text content. In real-world scenarios, noisy tables further hinder LMs’ symbolic reasoning capabilities. Consequently, while symbolic reasoning with numerical data is highly accurate, the generated code may be incorrect due to issues in program logic or data handling, leading to errors or unintended results. (Figure 1 - C4)
To address this, we first normalize the table structure and content, ensuring that each column follows a consistent format. We then propose a solution where LMs first engage in textual reasoning before generating symbolic reasoning programs. This preliminary textual reasoning step serves as a guide for subsequent symbolic reasoning, improving alignment with the task context. Our approach can be seen as encouraging LMs to think more thoroughly before reasoning, aligning with techniques like plan-and-solve [53]. By incorporating textual reasoning as a foundation, we enhance the accuracy and contextual relevance of symbolic reasoning. As demonstrated in Figure 2(c), this method achieves a higher accuracy of 59.1% for calculation-required questions. (Figure 1 - S4)
4 TableMaster: A Recipe for Table Understanding
@@
Based on findings in Section 3, we introduce a recipe and comprehensive framework, TableMaster, as shown in Figure 3. It integrates the propose solution proposed in Section 3 into a unified recipe for table understanding. The framework encompasses three key processes: Table Structure Understanding, Table Content Understanding, and Table Reasoning for QA. All notations are list at Appendix Q.
@@ 4.3 Table Content Understanding
Studies have shown that LMs can assess whether sufficient information is available to answer a question [54, 55]. We first prompt the LMs to estimate whether the constructed Table-of-Focus $\mathbb{T}^{F}_{a\times b}$, containing $C^{0}$, provides enough information to answer the given question $Q$. If not, additional column attributes from the candidate column set $\mathbf{C}$ are incrementally added from the ranked candidate headers until sufficient information is available or all relevant top headers have been utilized. Subsequently, a total of $a^{\prime}$ columns from $C$ are selected for further reasoning. We use re-construction to mitigate information loss during the table-of-focus construction process. The detailed re-construction algorithm can be found in Appendix J.
@@ 5 Experiments
Table 1: Performance comparison between TableMaster and previous work on WikiTQ and TabFact. The values in the table represent accuracy (%). The best result is bold, the second-best result is underlined, and the improvement over the previous best result is highlighted in green. ‘-’ indicates that the result values were not reported in the related papers. For all models in the table, results are obtained from a single inference run without any voting. Our method outperforms all other methods across both datasets and different language models.
Method WikiTQ TabFact
Text-to-SQL [56] 52.90 - - 64.71 - -
End-to-End QA [17] 51.84 - - 70.45 - -
Few-Shot QA [17] 52.56 - - 71.54 - -
Chain-of-Thought [17] 53.48 - - 65.37 - -
ReAcTable [43] 52.50 - - 74.40 - -
Binder [18] 56.74 58.86 50.51 79.17 84.63 78.16
Dater [16] 52.81 58.33 43.53 78.01 80.98 81.57
TabSQLify [42] 64.70 57.02 55.78 79.50 78.75 70.70
Chain-of-Table [17] 59.94 55.60 62.22 80.20 84.24 85.62
Tree-of-Table [45] 61.11 - - 81.92 - -
PoTable [41] - 64.73 65.56 - 88.93 87.06
@@ 5.1 Settings
We conduct extensive experiments to evaluate the performance of TableMaster. Specifically, we assess its effectiveness across three different table understanding datasets: WikiTQ [11] (table-based question answering), TabFact [12] (table-based fact verification), and FetaQA [57] (table-based free-form question answering). For WikiTQ and TabFact, following previous work [17, 46], we use exact match accuracy as the evaluation metric. For FetaQA, we evaluate performance using BLEU [58] and ROUGE [59] scores. We also conduct experiments on HiTab [60] and FinQA [61]. Tables are encoded in Markdown format before being input into language models, with or without addresses, depending on the specific case N.2.
Our experiments utilize OpenAI models hosted on Microsoft Azure. Unless otherwise stated, we set the temperature to 0 to ensure stable output while keeping all other hyperparameters at their default values. The models used in our evaluation include gpt-4o (gpt-4o-0806), gpt-4o-mini (gpt-4o-mini-0718), gpt-3.5-turbo (gpt-3.5-turbo-0125), o1 (o1-preview-0912), and o1-mini (o1-mini-0912). Additionally, we evaluate our methods on Llama-3.1-70B (Llama-3.1-70B-Instruct). The exact number of parameters for several LMs (e.g., GPT, o1) has not been publicly disclosed. Most parameter counts are estimates reported to provide context for understanding model performance. For more precise information, please refer to the original or future official documentation [62].
For comparison, we select several strong baselines, including both classic and state-of-the-art methods such as Binder [18], Dater [16], and Chain-of-Table [17]. Performance results for other methods not in this work are cited directly from their original or related papers, with sources indicated alongside the method names in the results table.
Further analysis and additional experiments on TableMaster can be found in the Appendix. The prompts used in TableMaster can be found in Appendix O, while other prompts used in this work are provided in Appendix P.